om/courses/24202/40122s/ Question 10 7 pts A motorcycle is moving at 16 m/s when its brakes are applied, bringing the cycle to rest in 3.5 s. To the nearest meter, how far does the motorcycle travel while coming to a stop?
Answer #1

initial velocity ,u =16m/s

final velocity ,v =0m/s

time =3.5s

First find deceleration

===================

Use formula v=u+at

0m/s 16m2/s a * 3.5s

0 16 + a * 3.5

a =-16/3.5 =-4.571m/s2

======================

Use Formula sut +1/2at2

s 16m/s 3.5s 1/2 4.571ms(3.5s)

16 3.5-0.5 4,571 3.52

ANSWER: s-28m

==================